# FullLoRA-AT: Efficiently Boosting the Robustness of Pretrained Vision Transformers

### v1

@@ Abstract
In recent years, the Vision Transformer (ViT) model has gradually become mainstream in various computer vision tasks, and the robustness of the model has received increasing attention. However, existing large models tend to prioritize performance during training, potentially neglecting the robustness, which may lead to serious security concerns. In this paper, we establish a new challenge: exploring how to use a small number of additional parameters for adversarial finetuning to quickly and effectively enhance the adversarial robustness of a standardly trained model. To address this challenge, we develop the novel LNLoRA module, incorporating a learnable layer normalization before the conventional LoRA module, which helps mitigate magnitude differences in parameters between the adversarial and standard training paradigms. Furthermore, we propose the FullLoRA-AT framework by integrating the learnable LNLoRA modules into all key components of ViT-based models while keeping the pretrained model frozen, which can significantly improve the model robustness via adversarial finetuning in a parameter-efficient manner. Extensive experiments on CIFAR-10, CIFAR-100, and Imagenette demonstrate the superiority of our proposed FullLoRA-AT framework. It achieves comparable robustness with full finetuning while only requiring about 5% of the learnable parameters. This also effectively addresses concerns regarding extra model storage space and enormous training time caused by adversarial finetuning.
1 Introduction
@@
Many existing research works have proposed various adversarial defense methods to enhance model robustness, including adversarial training [33, 62], adversarial purification [37], adversarial example detection [27], and certifiable defenses [8], among which adversarial training is the most commonly used and effective method. Adversarial training typically involves incorporating adversarial examples into the training set, enabling the model to resist such examples during training and significantly enhancing its robustness. Traditional approaches to adversarial training typically involve either training from scratch or fully finetuning from a pretrained model. This process requires updating the entire set of model parameters to develop a robust model.
@@
Figure 1: (a) A diagram of our proposed LNLoRA module. (b) A Vision Transformer (ViT) model utilizing our proposed FullLoRA-AT framework, which employs the LNLoRA module in the patch embedding layer, the multi-headed self-attention (MSA) block, the multi-layer perceptron (MLP) block and the MLP-based classification head. FullLoRA-AT enables the rapid enhancement of the adversarial robustness of the pretrained model through adversarial finetuning with a small number of trainable parameters.
@@
However, when dealing with ViT-based large models, the extensive number of parameters in these models and the immense size of training datasets make it challenging to apply adversarial training due to its considerable computational overhead. Many methods, such as FreeAT [41] and FastAT [56], have adopted a single-step generation of adversarial examples in place of the multi-step generation used in traditional adversarial training, significantly reducing the time required for adversarial training. Additionally, works like AGAT [57] have addressed the high computational cost of the attention mechanism in ViT models by adopting an attention-guided partial token discarding method to enhance the efficiency of adversarial training. Although the aforementioned methods have somewhat reduced the time required for adversarial training, they still necessitate updating all parameters of the model, resulting in significant additional model storage costs, while also incurring notable losses in robustness. Furthermore, in recent years, several methods have been proposed for the lightweight finetuning of pretrained large models in downstream tasks. Parameter-efficient language model tuning methods such as Adapters [18], Prefix-tuning [28], LoRA [20], UniPELT [35], AdaLoRA [63], and Compacter [34] have been introduced for efficient finetuning of large language models (LLMs) in downstream tasks. Recently, Aurora [52] has also been proposed for efficient downstream task finetuning of large-scale multimodal foundation models. These methods effectively enhance the performance of pretrained models in downstream tasks by finetuning a small number of learnable parameters. However, there has been no work focusing on designing a parameter-efficient approach to improve the adversarial robustness of pretrained models.
@@
To enhance the robustness of models pretrained using standard training in a lightweight manner, we draw inspiration from the LoRA method [20] and introduce the FullLoRA-AT framework. This framework aims to substantially improve the robustness of pretrained models against adversarial examples by adding only a small number of additional learnable parameters, which significantly reduces training time. Specifically, we first analyse the parameters from both adversarially and standardly trained models and reveal notable disparities in the parameters of scale and bias in the layernorm. These disparities can not be adequately addressed by the existing LoRA method. To tackle this issue, we introduce the novel LNLoRA module, which integrates a learnable layernorm layer within the existing LoRA module. Moreover, unlike the existing LoRA method which focuses only on the multi-headed self-attention (MSA) block and multi-layer perceptron (MLP) block in the Transformer layer, our approach introduces LNLoRA modules into more key components of the ViT, i.e., the patch embedding layer, MSA block, MLP block, and the classification head, as shown in Fig. 1. It is worth mentioning that our FullLoRA-AT method operates independently of existing advanced adversarial training techniques. This characteristic allows for seamless integration with others, leading to further improvement in the efficiency and performance of adversarial finetuning.
@@
We evaluate our FullLoRA-AT method on three datasets and several ViT-based models. Extensive experiments demonstrate that our proposed FullLoRA-AT method can significantly enhance the robustness of pretrained models against adversarial examples while substantially saving on additional training parameters and accelerating the training speed. In the case of ViT-B on the CIFAR10 dataset, compared to fully finetuned models, our method demonstrates remarkable efficiency, which requires only about 5% of the learnable parameters to achieve comparable robustness, i.e., incurring a mere 1%-2% loss in accuracy under the evaluation of various attacks, e.g., CW [3], PGD [33] and AutoAttack [9].
@@
1. We are the first to propose the task of rapidly enhancing the adversarial robustness of standardly pretrained models using a lightweight, parameter-efficient approach.
@@
3. We propose the FullLoRA-AT method, a lightweight parameter-efficient approach to rapidly improve the robustness of the pretrained model, which significantly reduces the number of learnable parameters, thereby substantially saving on model storage space and reducing the time required for model finetuning.
2 Related Work
2.1 Vision Transformer
@@
Vision Transformer (ViT), since its introduction by [13], has revolutionized the field of computer vision, marking a significant shift from the conventional Convolutional Neural Networks (CNNs) [17, 47, 21] that dominated the field for years. ViT employs self-attention mechanisms, typical of transformer models in natural language processing [50], to process images, which distinguishes it from the localized and hierarchical nature of CNNs [25, 42]. In image classification, ViT [13] has shown remarkable success, achieving state-of-the-art performance on benchmarks like ImageNet [40]. Later, various variants of ViT have been proposed to improve the efficiency and performance of ViT models, i.e., DeiT [48], Swin [31], MetaFormer [61] and PSLT [58]. With the great success of ViT in image classification, the application of ViT-based models has also extended to other visual tasks, such as object detection [2, 14, 53, 55], semantic segmentation [65, 22, 44, 7], image super resolution [60, 29, 64]. In the various visual tasks mentioned above, ViT-based models have achieved significant performance improvements and have gradually become mainstream models in the field of computer vision.
2.2 Adversarial Training
@@
Existing research has proposed numerous methods for the adversarial defense of models, such as adversarial training [33], adversarial purification [37], adversarial example detection [27], and certifiable defense [8], with adversarial training being one of the most effective. Adversarial training is first introduced by PGD [33], which improves model accuracy under adversarial attacks by incorporating generated adversarial examples into the training set. Researches like TRADES [62] and MART [54] further enhance model robustness by optimizing the objective function during adversarial training. Due to the significant time required to generate adversarial examples in adversarial training, methods like FreeAT [41] and FastAT [56] have markedly accelerated this process through single-step attacks while maintaining high model robustness. For adversarial training on ViT models, recent studies [10, 43] have improved this process from the perspectives of model structure, training strategy, and training tricks. Mo et al. [36] improves the adversarial robustness of ViTs by randomly masking gradients from some attention blocks and masking perturbations on some patches during training. AGAT [57] addresses the computational expense of the attention mechanism in ViT models by discarding some tokens guided by attention, thus saving time in adversarial training. Rebuffi et al. [38] trains separate classification tokens of a ViT for the clean and adversarial domains, achieving good clean and robust accuracy, which can be further balanced through a model soup approach. AutoLoRA [59] uses the LoRA [20] mechanism to decouple features of clean and adversarial examples, thereby enhancing the robustness of pretrained large models in downstream tasks after finetuning.
@@
Most existing adversarial training methods involve training or finetuning the entire set of model parameters. In contrast, our work freezes most of the pretrained model parameters and proposes to enhance the robustness of pretrained models in a lightweight manner through a small number of additional training parameters.
2.3 Parameter-efficient Model Finetune
@@
Large models pretrained on extensive datasets exhibit commendable generalizability, making finetuning on such models for downstream tasks a prevalent paradigm. In recent years, numerous parameter-efficient finetuning methods have been proposed for efficiently optimizing models for downstream tasks. In natural language processing (NLP), methods such as Adapters [18], Prefix-tuning [28], LoRA [20], UniPELT [35], AdaLoRA [63] and Compacter [34] have been introduced. They add a small number of additional learnable parameters to the pretrained model to facilitate rapid and efficient finetuning for downstream NLP tasks. In visual tasks, Visual Prompt [23] introduces only a small number of trainable parameters in the input space while keeping the model backbone frozen, often outperforming full finetuning. Furthermore, with the recent rise of multimodal models, many works have focused on the parameter-efficient finetuning of these models. VL-Adapter [45] introduces adapter-based parameter-efficient transfer learning techniques to vision-language models, and Aurora [52] applies CP decomposition to model parameters and proposes the informative context enhancement and gated query transformation module for lightweight finetuning.
@@
3 Method
In this section, we first provide a brief overview of the Vision Transformer [13] (ViT) model, which has been widely used in various computer vision tasks, as well as the existing Low-Rank Adaptation (LoRA) method [20], known for its efficiency in lightweight finetuning of pretrained models for various downstream tasks. Following this, we introduce our novel FullLoRA-AT method, a lightweight approach designed to enhance the adversarial robustness of standardly pretrained models in a parameter-efficient manner.
3.1 Preliminary
@@
3.1.1 Vision Transformer
In recent years, the Vision Transformer (ViT) has achieved remarkable performance improvements in various visual tasks and has gradually become the mainstream model for these tasks.
The ViT model first resizes the input image $x\in\mathbb{R}^{H\times W\times C}$ into different patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ is the resolution of the image, $C$ is the number of channels in the image, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the number of patches after resizing. ViT initially maps these image patches to features through a patch embedding layer $E\in\mathbb{R}^{(P^{2}\cdot C)\times D}$ , concatenates them with a special classification token $x_{class}$, and adds position embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ to obtain the embedded patches $z_{0}$. These embedded patches are then transformed by $L$ consecutive Transformer layers, each consisting of a multi-headed self-attention (MSA) and multi-layer perceptron (MLP) block, resulting in the output $z_{L}$ of the entire Transformer encoder. Finally, the feature $z_{L}^{0}$ at the classification token position of $z_{L}$ is processed through an MLP-based classification head layer to produce the model’s final classification result $y$. The overall process can be formulated as follows:
$z_{0}=[x_{\text{class}};x_{p}^{1}E;x_{p}^{2}E;\cdots;x_{p}^{N}E]+E_{pos},$ (1)
@@ 3.1.1 Vision Transformer
where $\mathrm{LN}$ is the layernorm, $\mathrm{MSA}$ is the multi-headed self-attention block, $\mathrm{MLP}$ is the multi-layer perceptron block and $\mathrm{MLPHead}$ is the MLP-based classification head.
3.1.2 Low-Rank Adaptation
Low-rank adaptation (LoRA) is an efficient method for finetuning pretrained models on downstream tasks. It utilizes a low-rank approach for updating additional model weights rather than directly finetuning the pretrained model, effectively reducing the number of parameters and training costs.
@@ 3.1.2 Low-Rank Adaptation
Figure 2: Visualization of the average absolute values of different parameters across the 12 Transformer layers of the ViT model. (a) presents the model with standard training, (b) presents the model with adversarial training, and (c) presents the differences in model parameters between (a) and (b).
3.2 FullLoRA Adversarial Training
Existing Vision Transformer (ViT) models have achieved promising results in various visual tasks. However, during model training, there is often a greater focus on performance, potentially leading to a neglect of the model’s robustness. In response to this situation, we propose a new task setting: for pretrained models obtained through standard training, we aim to enhance the adversarial robustness quickly using a parameter-efficient approach for finetuning. To this end, we introduce FullLoRA Adversarial Training, a method that minimally increases parameters to enhance model robustness. Compared to the commonly used full finetuning approach, our FullLoRA-AT introduces a significantly smaller number of additional training parameters, thereby saving on both training time and model storage costs during finetuning.
In the following, we will provide a detailed introduction to our proposed FullLoRA-AT method.
3.2.1 LNLoRA

### v2

# FullLoRA: Efficiently Boosting the Robustness of Pretrained Vision Transformers

@@ Abstract
In recent years, the Vision Transformer (ViT) model has gradually become mainstream in various computer vision tasks, and the robustness of the model has received increasing attention. However, existing large models tend to prioritize performance during training, potentially neglecting the robustness, which may lead to serious security concerns. In this paper, we establish a new challenge: exploring how to use a small number of additional parameters for adversarial finetuning to quickly and effectively enhance the adversarial robustness of a standardly trained model. To address this challenge, we develop novel LNLoRA module, incorporating a learnable layer normalization before the conventional LoRA module, which helps mitigate magnitude differences in parameters between the adversarial and standard training paradigms. Furthermore, we propose the FullLoRA framework by integrating the learnable LNLoRA modules into all key components of ViT-based models while keeping the pretrained model frozen, which can significantly improve the model robustness via adversarial finetuning in a parameter-efficient manner. Extensive experiments on several datasets demonstrate the superiority of our proposed FullLoRA framework. It achieves comparable robustness with full finetuning while only requiring about 5% of the learnable parameters. This also effectively addresses concerns regarding extra model storage space and enormous training time caused by adversarial finetuning.
Index Terms:
Adversarial Training, Parameter-Efficient, Pretrained Model
I Introduction
@@
Many existing research works have proposed various adversarial defense methods to enhance model robustness, including adversarial training [22, 23, 24, 25, 26], adversarial purification [27], adversarial example detection [28], and certifiable defenses [29], among which adversarial training is the most commonly used and effective method. Adversarial training typically involves incorporating adversarial examples into the training set, enabling the model to resist such examples during training and significantly enhancing its robustness. Traditional approaches to adversarial training typically involve either training from scratch or fully finetuning from a pretrained model. This process requires updating the entire set of model parameters to develop a robust model.
@@
Figure 1: (a) A diagram of our proposed LNLoRA module. (b) A Vision Transformer (ViT) model utilizing our proposed FullLoRA framework, which employs the LNLoRA module in the patch embedding layer, the multi-headed self-attention (MSA) block, the multi-layer perceptron (MLP) block and the MLP-based classification head. FullLoRA enables the rapid enhancement of the adversarial robustness of the pretrained model through adversarial finetuning with a small number of trainable parameters.
@@
However, when dealing with ViT-based large models, the extensive number of parameters in these models and the immense size of training datasets make it challenging to apply adversarial training due to its considerable computational overhead. Many methods, such as FreeAT [30] and FastAT [31], have adopted a single-step generation of adversarial examples in place of the multi-step generation used in traditional adversarial training, significantly reducing the time required for adversarial training. Additionally, works like AGAT [32] have addressed the high computational cost of the attention mechanism in ViT models by adopting an attention-guided partial token discarding method to enhance the efficiency of adversarial training. Although the aforementioned methods have somewhat reduced the time required for adversarial training, they still necessitate updating all parameters of the model, resulting in significant additional model storage costs, while also incurring notable losses in robustness. Furthermore, in recent years, several methods have been proposed for the lightweight finetuning of pretrained large models in downstream tasks. Parameter-efficient language model tuning methods such as Adapters [33], Prefix-tuning [34], LoRA [35], UniPELT [36], AdaLoRA [37], and Compacter [38] have been introduced for efficient finetuning of large language models (LLMs) in downstream tasks. Recently, Aurora [39] and VeRA [40] have also been proposed for efficient downstream task finetuning of large-scale multimodal foundation models. These methods effectively enhance the performance of pretrained models in downstream tasks by finetuning a small number of learnable parameters. However, there has been few work focusing on designing a parameter-efficient approach to improve the adversarial robustness of pretrained models.
@@
To enhance the robustness of models pretrained using standard training in a lightweight manner, we draw inspiration from the LoRA method [35] and introduce the FullLoRA framework. This framework aims to substantially improve the robustness of pretrained models against adversarial examples by adding only a small number of additional learnable parameters, which significantly reduces training time. Specifically, we first analyse the parameters from both adversarially and standardly trained models and reveal notable disparities in the parameters of scale and bias in the layernorm. These disparities can not be adequately addressed by the existing LoRA method, which is demonstrated in Fig. 3 of Section IV-G. To tackle this issue, we introduce the novel LNLoRA module, which integrates a learnable layernorm layer within the existing LoRA module. Moreover, unlike the existing LoRA method which focuses only on the multi-headed self-attention (MSA) block and multi-layer perceptron (MLP) block in the Transformer layer, our approach introduces LNLoRA modules into more key components of the ViT, i.e., the patch embedding layer, MSA block, MLP block, and the classification head, as shown in Fig. 1. It is worth mentioning that our FullLoRA method is independent of existing advanced adversarial training techniques. This characteristic allows for seamless integration with other methods, leading to further improvement in the efficiency and effectiveness of adversarial finetuning.
@@
We evaluate our FullLoRA method on several datasets and ViT-based models. Extensive experiments demonstrate that our proposed FullLoRA method can significantly enhance the robustness of pretrained models against adversarial examples while substantially saving on additional training parameters and accelerating the training speed. In the case of ViT-B on the CIFAR10 dataset, compared to fully finetuned models, our method demonstrates remarkable efficiency, which requires only about 5% of the learnable parameters to achieve comparable robustness, i.e., incurring a mere 1%-2% loss in accuracy under the evaluation of various attacks, e.g., CW [41], PGD [22] and AutoAttack [42].
@@
1. We propose the task of enhancing the adversarial robustness of standardly pretrained models using a lightweight, parameter-efficient approach.
@@
3. We propose the FullLoRA method, a lightweight parameter-efficient approach to rapidly improve the robustness of the pretrained model, which significantly reduces the number of learnable parameters, thereby substantially saving on model storage space and reducing the time required for model finetuning.
II Related Work
II-A Vision Transformer
@@
Vision Transformer (ViT), since its introduction by [10], has revolutionized the field of computer vision, marking a significant shift from the conventional Convolutional Neural Networks (CNNs) [2, 3, 43] that dominated the field for years. ViT employs self-attention mechanisms, typical of transformer models in natural language processing [44], to process images, which distinguishes it from the localized and hierarchical nature of CNNs [45, 1]. In image classification, ViT [10] has shown remarkable success, achieving state-of-the-art performance on benchmarks like ImageNet [46]. Later, various variants of ViT have been proposed to improve the efficiency and performance of ViT models, i.e.., DeiT [11], Swin [12], MetaFormer [47] and PSLT [48]. With the great success of ViT in image classification, the application of ViT-based models has also extended to other visual tasks, such as object detection [13, 49, 50, 51], semantic segmentation [14, 52, 53, 54], image super resolution [55, 56, 57]. In the various visual tasks mentioned above, ViT-based models have achieved significant performance improvements and have gradually become mainstream models in the field of computer vision.
II-B Adversarial Training
@@
Existing research has proposed numerous methods for the adversarial defense of models, such as adversarial training [22], adversarial purification [27], adversarial example detection [28], and certifiable defense [29], with adversarial training being one of the most effective. Adversarial training is first introduced by [22], which improves model accuracy under adversarial attacks by incorporating generated adversarial examples into the training set. Researches like TRADES [23] and MART [58] further enhance model robustness by optimizing the objective function during adversarial training. RiFT [24] proposes to exploit the redundant capacity for robustness by fine-tuning the adversarially trained model on its non-robust-critical module to en-hance generalization without compromising adversarial robustness. TWINS [25] explores how to maintain the robustness in the pre-trained model when learning the downstream task, and proposes a novel statistics-based framework, which is effective on a wide range of image classification datasets in terms of both generalization and robustness. PMG-AFT [26] leverages supervi-sion from the original pre-trained model by carefully designing an auxiliary branch, to enhance the model’s zero-shot adversarial robustness. Due to the significant time required to generate adversarial examples in adversarial training, methods like FreeAT [30] and FastAT [31] have markedly accelerated this process through single-step attacks while maintaining high model robustness. For adversarial training on ViT models, recent studies [59, 60] have improved this process from the perspectives of model structure, training strategy, and training tricks. Mo et al. [61] improves the adversarial robustness of ViTs by randomly masking gradients from some attention blocks and masking perturbations on some patches during training. AGAT [32] addresses the computational expense of the attention mechanism in ViT models by discarding some tokens guided by attention, thus saving time in adversarial training. Rebuffi et al. [62] trains separate classification tokens of a ViT for the clean and adversarial domains, achieving good clean and robust accuracy, which can be further balanced through a model soup approach. AutoLoRA [63] uses the LoRA [35] mechanism to decouple features of clean and adversarial examples, thereby enhancing the robustness of pretrained models in downstream tasks after finetuning.
@@
Most existing adversarial training methods involve training or finetuning the entire set of model parameters. In contrast, our work freezes most of the pretrained model parameters and proposes to enhance the robustness of pretrained models in a lightweight manner through a small number of additional trainable parameters.
II-C Parameter-efficient Model Finetune
@@
Large models pretrained on extensive datasets exhibit commendable generalizability, making finetuning on such models for downstream tasks a prevalent paradigm. In recent years, numerous parameter-efficient finetuning methods have been proposed for efficiently optimizing models for downstream tasks. In natural language processing (NLP), methods such as Adapters [33], Prefix-tuning [34], LoRA [35], UniPELT [36], AdaLoRA [37] and Compacter [38] have been introduced. They add a small number of additional learnable parameters to the pretrained model to facilitate rapid and efficient finetuning for downstream NLP tasks. In visual tasks, Visual Prompt [64] introduces only a small number of trainable parameters in the input space while keeping the model backbone frozen, often outperforming full finetuning. Furthermore, with the recent rise of multimodal models, many works have focused on the parameter-efficient finetuning of these models. VL-Adapter [65] introduces adapter-based parameter-efficient transfer learning techniques to vision-language models, and Aurora [39] applies CP decomposition to model parameters and proposes the informative context enhancement and gated query transformation module for lightweight finetuning. Recently, VeRA [40] proposes to use a single pair of low-rank matrices shared across all layers and learn small scaling vectors instead, which shows the effectiveness both on language and vision tasks.
@@
III Method
In this section, we first provide a brief overview of the Vision Transformer [10] (ViT) model, which has been widely used in various computer vision tasks, as well as the existing Low-Rank Adaptation (LoRA) method [35], known for its efficiency in lightweight finetuning of pretrained models for various downstream tasks. Following this, we introduce our novel FullLoRA method, a lightweight approach designed to enhance the adversarial robustness of standardly pretrained models in a parameter-efficient manner.
III-A Preliminary
@@
Vision Transformer. In recent years, the Vision Transformer (ViT) has achieved remarkable performance improvements in various visual tasks and has gradually become the mainstream model for these tasks.
The ViT model first resizes the input image $x\in\mathbb{R}^{H\times W\times C}$ into different patches $x_{p}\in\mathbb{R}^{N\times(P^{2}\cdot C)}$, where $(H,W)$ is the resolution of the image, $C$ is the number of channels in the image, $(P,P)$ is the resolution of each image patch, and $N=HW/P^{2}$ is the number of patches after resizing. ViT initially maps these image patches to features through a patch embedding layer $E\in\mathbb{R}^{(P^{2}\cdot C)\times D}$, concatenates them with a special classification token $x_{class}$, and adds position embedding $E_{pos}\in\mathbb{R}^{(N+1)\times D}$ to obtain the embedded patches $z_{0}$, as:
$z_{0}=[x_{\text{class}};x_{p}^{1}E;x_{p}^{2}E;\cdots;x_{p}^{N}E]+E_{pos}.$ (1)
These embedded patches are then transformed by $L$ consecutive Transformer layers, each consisting of a multi-headed self-attention (MSA) and multi-layer perceptron (MLP) block, resulting in the output $z_{L}$ of the entire Transformer encoder, which can be formulated as:
@@
where $\mathrm{LN}$ is the layernorm, $\mathrm{MSA}$ is the multi-headed self-attention block, $\mathrm{MLP}$ is the multi-layer perceptron block. Finally, the feature $z_{L}^{0}$ at the classification token position of $z_{L}$ is processed through an MLP-based classification head layer to produce the model’s final classification result $y$:
@@
where $\mathrm{MLPHead}$ is the MLP-based classification head.
Low-Rank Adaptation. Low-rank adaptation (LoRA) is an efficient method for finetuning pretrained models on downstream tasks. It utilizes a low-rank approach for updating additional model weights rather than directly finetuning the pretrained model, effectively reducing the number of parameters and training costs.
@@
III-B FullLoRA Adversarial Training
Existing Vision Transformer (ViT) models have achieved promising results in various visual tasks. However, during model training, there is often a greater focus on performance, potentially leading to a neglect of the model’s robustness. In response to this situation, we propose a new task setting: for pretrained models obtained through standard training, we aim to enhance the adversarial robustness quickly using a parameter-efficient approach for finetuning. To this end, we introduce FullLoRA Adversarial Training, a method which enhances model robustness with minimal additional training parameters. Compared to the commonly used full finetuning approach, our FullLoRA introduces a significantly smaller number of additional trainable parameters, thereby saving on both training time and model storage costs during finetuning.
In the following, we will provide a detailed introduction to our proposed FullLoRA method.
@@ III-B FullLoRA Adversarial Training
Figure 2: Visualization of the average absolute values of different parameters across the 12 Transformer layers of the ViT model. (a) presents the model with standard training, (b) presents the model with adversarial training, and (c) presents the differences in model parameters between (a) and (b). The experiment is repeated by 5 times, the solid line represents the mean value of the 5 experiments, while the shaded area represents the standard deviation.
III-B1 LNLoRA